In how many ways can 12 books be divided among 4 students so that each student receives 3 books?
1
Expert's answer
2020-03-02T09:55:55-0500

Number of ways to choose 3 books for the first student out of 12:C312

Number of ways to choose 3 books for the second student out of 9:C39

Number of ways to choose 3 books for the third student out of 6:C36

Number of ways to choose 3 books for the fourth student out of 3:1

Total options:

"C^3_{12}\cdot C^3_9\cdot C^3_6\cdot 1=\frac{12!\cdot9!\cdot6!}{3!\cdot9!\cdot3!\cdot6!\cdot3!\cdot3!}=369600"